# Velocities of Free Floaters in a Sea of Stars

### v1

@@ I Introduction
$v_{c}\simeq\Bigg({Gm_{p}\over 0.12a_{p}}\Bigg)^{1/2}{m_{p}\over m_{p}+m_{i}}=(2.7\,\mathrm{km\,s}^{-1})\Bigg({m_{p}\over m_{J}}\Bigg)^{1/2}\Bigg({a_{p}\over 1\mathrm{AU}}\Bigg)^{-1/2}{m_{p}\over m_{p}+m_{i}}$
@@
$v_{c}\simeq\Bigg({Gm_{0}\over 0.12a_{p}}\Bigg)^{1/2}{m_{0}\over m_{0}+m_{1}}=(87\,\mathrm{km\,s}^{-1})\Bigg({m_{0}\over m_{\odot}}\Bigg)^{1/2}\Bigg({a_{p}\over 1\mathrm{AU}}\Bigg)^{-1/2}{m_{0}\over m_{0}+m_{1}}$
@@ II Gravitational Scattering
${\partial\left(\tilde{v}^{2}\right)\over\partial\tilde{t}}={m_{p}\over m}\Bigg[-{\sqrt{\pi}\over 2\tilde{v}}\mathrm{erf}(\tilde{v})+\left(1+{m\over m_{p}}\right)\exp\left({-\tilde{v}^{2}}\right)\Bigg].$ (8)

### v2

@@ I Introduction
$v_{c}$ $\simeq\Bigg({Gm_{p}\over 0.12a_{p}}\Bigg)^{1/2}{m_{p}\over m_{p}+m_{i}}$
[…]
@@
$v_{c}$ $\simeq\Bigg({Gm_{0}\over 0.12a_{p}}\Bigg)^{1/2}{m_{0}\over m_{0}+m_{1}}$
[…]
@@ II Gravitational Scattering
$\times\left({\sigma\over 1\,\mathrm{km\,s}^{-1}}\right)^{3}\left({m\over\,{\rm M}_{\odot}}\right)^{-2}.$
@@
${\partial\tilde{v}^{2}\over\partial\tilde{t}}={m_{p}\over m}\Bigg[-{\sqrt{\pi}\over 2\tilde{v}}\mathrm{erf}(\tilde{v})+\left(1+{m\over m_{p}}\right)\exp\left({-\tilde{v}^{2}}\right)\Bigg].$ (8)
@@ IV.1 Evolution of Velocity Distribution
$\bar{F}({\tilde{v}},{\tilde{T}})={{\tilde{v}}^{2}\over{\tilde{T}}}\left[{\sqrt{\pi}\over 2{\tilde{v}}}\exp({\tilde{v}}^{2})\,{\rm erf}({\tilde{v}})-1\right]$ (25)